# Revisiting XL-MIMO Channel Estimation: When Dual-Wideband Effects Meet Near Field

### v2

@@ III-B Angular-Delay Representation of SnS XL-MIMO Channel
$\lim_{N,K\to\infty}[\mathbf{X}]_{n,k}=\begin{cases}\text{nonzeros},&(n,k)\in\mathcal{A}_{2},\\ 0,&(n,k)\notin\mathcal{A}_{2},\end{cases}$ (16)
where $\mathcal{A}_{2}=\cup_{l,p}\mathcal{A}_{l,2,p}$ with
$\mathcal{A}_{l,2,p}$ =$=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l}+N\epsilon_{l,p},N),\right.$ (17)
$k$ $\left.=\mathrm{mod}(k_{l}+f_{s}\tau_{l},K),\forall(n_{l},k_{l})\in\mathcal{A}_{l,1}\right\},$
and $\mathrm{mod}(a,m)$ is the modulus of $a$ for $m$. In addition, $\epsilon_{l,p}=\psi_{l}-2n_{l,e}\varphi_{l}+\eta_{l}(p-1)$ with $\eta_{l}={2N\varphi_{l}}/{P}$ and $p\in\{p\in\mathbb{N}^{+}\mid\psi_{l}-2n_{l,e}\varphi_{l}\leq\eta_{l}p\leq\psi_{l}-2n_{l,s}\varphi_{l}\}$. $\mathcal{A}_{l,1}\triangleq\left\{(n_{l},k_{l})\in\mathbb{Z}^{2}\mid I_{l,e}\leq n_{l}\leq I_{l,s},J_{l,e}\leq k_{l}\leq J_{l,s}\right\}$ with $I_{l,e}=\min_{k}\lceil(i_{l,e}(k)+1)N/2\rceil$, $I_{l,s}=\max_{k}\lceil(i_{l,s}(k)+1)N/2\rceil$, $J_{l,e}=\min_{n}j_{l,e}(n)$, and $J_{l,s}\max_{n}j_{l,s}(n)$, $\forall n\in\boldsymbol{\phi}_{l}$, where
$i_{l,e}(k)$ $=2(a_{k}-b_{k}n_{l,e}),\quad i_{l,s}(k)=2(a_{k}-b_{k}n_{l,s}),$ (18)
$j_{l,e}(n)$ $=(t_{l,n}-{2}/K)K,\quad j_{l,s}(n)=(t_{l,n}+{2}/K)K,$ (19)
with $a_{k}={f_{k}\psi_{l}}/{f_{c}}$, $b_{k}={2f_{k}\varphi_{l}}/{f_{c}}$, and $t_{l,n}=(n{f_{s}\psi_{l}}-n^{2}{f_{s}\varphi_{l}})/(Kf_{c})$.
From (17), the non-zero region of $\mathbf{X}$ is synthesized from the shifted non-zero regions of $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\psi_{l},\varphi_{l})\mathbf{F}_{\mathrm{D}}$ across angular and delay domains for all $l$. The shift distance are jointly determined by $N\epsilon_{l,p}$ and $f_{s}\tau_{l}$. In addition, The index of significant angular-delay components $(n,k)$ within $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ form a parabolic shape, which arises because both $j_{l,e}(k)$ and $j_{l,s}(k)$ are quadratic functions of $n$.
Proof: Please see Appendix A.
Moreover, we analyze the effects of the spherical wavefront, SnS properties, and dual-wideband effects on the sparsity structure of $\mathbf{X}$ based on Lemma 1 from the perspectives of the angular and delay spread. 1) Angular Spread: The spherical wavefront introduces angular spread of $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})$ in angular domain, ranging from $a_{k}-b_{k}n_{l,e}$ to $a_{k}-b_{k}n_{l,s}$ due to the inclusion of a quadratic term, $\varphi_{l}n^{2}$. Meanwhile, it affects the variations of $\epsilon_{l,p}$. The SnS properties influence the values of $n_{l,e}$ and $n_{l,s}$, thereby affecting the extent of the angular spread. Additionally, the frequency-wideband effect induces variations in $i_{l,e}(k)$ and $i_{l,s}(k)$ across subcarriers $k$, which further broadens the angular dispersion across the system bandwidth. 2) Delay Spread: The spatial-wideband effect causes variations in $j_{l,e}(n)$ and $j_{l,s}(n)$ across antennas $n$, leading to delay spread. The SnS properties define the range of variation for $n$, while the spherical wavefront imposes a quadratic trend on the delay spread as a function of $n$.
@@
In summary, the spherical wavefront effect, SnS properties, and dual-wideband effects collectively influence the angular and delay spreads, reshaping the sparse patterns of SnS dual-wideband channels in the angular-delay domain, i.e.,
@@
  • Global Block Sparsity: The angular-delay channel contains $L$ non-zero block corresponding to $L$ propagation paths. The index range of each non-zero block is determined by $\mathcal{A}_{l,2}$.
  • Local Common-Delay Sparsity: Due to the spherical wavefront effect, within $\mathcal{A}_{2}$, different delay indices correspond to varying angular spreads, which occur in clusters and exhibit high correlation.
@@
To enhance intuitive understanding of the SnS XL-MIMO channel, we provide an example with the following parameters: $N=K=1024$, $f_{s}=3$GHz, $f_{c}=30$GHz, $r_{l}=10$m, $\vartheta_{l}=\pi/5$, and $\boldsymbol{\phi}_{l}\in[64,959]$. Fig. 2 illustrates the block sparsity of the SnS XL-MIMO channel in the angular-delay domain. Given the specified parameters, we have $i_{l,e}(0)=0.64$, $i_{l,s}(0)\approx 0.8$, $j_{l,e}(64)=513$, and $j_{l,s}(64)\approx 517$. Similarly, $i_{l,e}(1023)\approx 0.88$, $i_{l,s}(1023)\approx 0.72$, $j_{l,e}(959)=545$, and $j_{l,s}(959)\approx 549$. As depicted in Fig. 2(a) and Fig. 2(b), these values provide accurate boundaries for the ranges of angular and delay spread. Additionally, the figures show that both the angular and delay spectra shift to the right due to $\cos\vartheta_{l}>0$ and the monotonicity of $t_{n}$ with respect to $n$.
Furthermore, we obtain the following values: $I_{l,e}=841$, $I_{l,s}=962$, $J_{l,e}=513$, and $J_{l,s}=549$. Fig. 2(c) illustrates the non-zero region of $\mathbf{F}^{\mathrm{A}}\boldsymbol{\Theta}(\psi_{l},\varphi_{l})\mathbf{F}^{*}_{\mathrm{D}}$, which is accurately constrained to $\mathcal{A}_{l,1}$, validating the effectiveness of Lemma 1. Additionally, we have $f_{s}\tau_{l}=100$, $\epsilon_{l,1}=\psi_{l}-2n_{l,s}\varphi_{l}\approx 0.4$, and $\epsilon_{l,P}=\psi_{l}-2n_{l,e}\varphi_{l}=0.32$, corresponding to $N\epsilon_{l,1}=410$ and $N\epsilon_{l,P}=328$, respectively. Consequently, Fig. 2(d) and Fig. 2(e) represent shifted versions of Fig. 2(c). The shift distances in the angular domain are $696$ and $614$, respectively, while the shift distance in the delay domain is $100$. Finally, by concentrating the shifted versions across all $\epsilon_{l,p}$, the sparse structure depicted in Fig. 2(f) is obtained. It can be observed that the angular range of $\mathcal{A}_{l,2}$ is broader than that of $\mathcal{A}_{l,1}$ due to the composite effect.
(a) Angular spread for $k=0$ and $K-1$.
(b) Delay spread for $n=0$ and $N_{\mathrm{R}}-1$.
(c) $\mathcal{A}_{l,1}$.
(d) $\mathcal{A}_{l,p,1}$ with $p=1$.
(e) $\mathcal{A}_{l,p,1}$ with $p=P$
(f) $\mathcal{A}_{l,2}$.
@@
Compared to conventional massive MIMO-OFDM channels, the proposed SnS dual-wideband XL-MIMO channel exhibits distinct behavior. Specifically, in addition to beam squint, the spherical wavefront effect and SnS properties contribute to angular spread, resulting in broader angular diffusion. Moreover, the quadratic phase variations introduced by the spherical wavefront reshape the trend of delay variations across antennas. Notably, Lemma 1 is applicable to conventional MIMO-OFDM channels. When the spherical wavefront simplifies to a plane wavefront and SnS properties are disregarded, both $\varphi_{l}$ and $\kappa_{l,n}$ can be omitted. Thus, (18) simplifies to $i_{l,e}(k)=i_{l,s}(k)$ and $t_{l,n}=nf_{s}\psi_{l}$. Consequently, the significant components are uniformly distributed within the non-zero region, as illustrated in [6].
IV SnS Dual-Wideband Channel Estimation for XL-MIMO Systems
In this section, leveraging the sparsity outlined in Lemma 1, we first formulate the SnS dual-wideband XL-MIMO channel estimation problem as a sparse signal recovery task. Subsequently, to account for the global block sparsity and local common-delay sparsity, we introduce a novel column-wise sparse prior model tailored to this structure.
@@ IV-A Problem Formulation
Motivated by the sparsity of SnS dual-wideband XL-MIMO channels in the angular-delay domain, the channel estimation problem can be formulated as a multiple measurement vectors (MMV) based sparse recovery problem. Utilizing the angular-delay representation in (15), the received signal model in (3) can be further written as
$\mathbf{Y}=\mathbf{W}\mathbf{F}_{\mathrm{A}}\mathbf{X}\mathbf{F}_{\mathrm{D}}+{\mathbf{N}}=\boldsymbol{\Psi}\mathbf{X}\mathbf{F}_{\mathrm{D}}+{\mathbf{N}},$ (20)
@@
$\tilde{\mathbf{Y}}={\mathbf{Y}}\mathbf{F}_{\mathrm{D}}^{\mathrm{H}}=\boldsymbol{\Psi}\mathbf{X}+\tilde{\mathbf{N}},$ (21)
where $\tilde{\mathbf{Y}}={\mathbf{Y}}\mathbf{F}_{\mathrm{D}}^{\mathrm{H}}\in\mathbb{C}^{M\times K}$ and $\tilde{\mathbf{N}}={\mathbf{N}}\mathbf{F}_{\mathrm{D}}^{\mathrm{H}}\in\mathbb{C}^{M\times K}$ denotes the equivalent received pilot signal and noise matrix.
In this paper, our primary objective is to develop an effective estimation algorithm for accurately reconstructing $\mathbf{X}$ or $\mathbf{H}$ according to $\mathbf{Y}$ and $\boldsymbol{\Psi}$. In the existing literature, several sparse signal recovery algorithms have been employed to address the problem in (21), including orthogonal matching pursuit (OMP) [28, 29] and optimization-based methods [30, 31]. However, these methods fail to effectively exploit the inherent sparsity structure of the angular-delay channel and often require additional parameters, such as the rank of $\mathbf{X}$ or the number of multipaths, which may not always be available.
Considering the block sparsity of $\mathbf{X}$, Bayesian inference techniques such as approximate message passing (AMP) and sparse Bayesian learning (SBL) have shown superior recovery performance in channel estimation tasks [32, 33, 27, 34]. However, the effectiveness of these Bayesian methods relies heavily on the accuracy of the prior model. To address this, we propose a novel column-wise hierarchical sparse prior model, as detailed in the next section.

### v3

@@ III-B Angular-Delay Representation of SnS XL-MIMO Channel
$[\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}]_{n,k}=\begin{cases}\text{nonzeros},&(n,k)\in\mathcal{A}_{l},\\ 0,&(n,k)\notin\mathcal{A}_{l},\end{cases}$ (24)
where the set $\mathcal{A}_{l}$ is defined as
$\mathcal{A}_{l}\triangleq\left\{(n,k)\in\mathbb{Z}^{2}\mid I_{l,e}\leq n\leq I_{l,s},J_{l,e}\leq k\leq J_{l,s}\right\},$ (25)
with
$I_{l,e}$ $=\min_{k}\lceil(i_{l,e}(k)+1)N/2\rceil,$ (26)
$I_{l,s}$ $=\max_{k}\lceil(i_{l,s}(k)+1)N/2\rceil,$
$J_{l,e}$ $=\max_{n}(t_{l,n}+{2}/{K}+1/2)K,$
$J_{l,s}$ $=\min_{n}(t_{l,n}-{2}/{K}+1/2)K,$
indicating the boundaries.
In addition, within the region $\mathcal{A}_{l}$, each column corresponds to the angular response associated with a specific propagation delay. The width of the significant spatial frequency components is primarily determined by the frequency-width effect. Assuming the starting spatial frequency of the $k$-th column is denoted by $i_{l,k}\in[i_{l,e}^{\min},i_{l,s}^{\max}]$, the corresponding angular spread can be approximated as
$B_{l,k}=\frac{i_{l,k}f_{s}}{f_{c}}N.$ (27)
As a result, the significant components of $\boldsymbol{\Xi}_{l}$ within $\mathcal{A}_{l}$ are confined to narrow bands along each column, thereby exhibiting a pronounced column-wise clustered sparsity pattern.
To summarize, by leveraging (24)-(27), the sparse structure of $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{*}$ can be accurately characterized. In particular, equation (24) delineates the overall boundaries of the nonzero region, while (27) provides a more precise description of the angular spread in each column, further refining the support of the dominant components.
Proof.
Please see Appendix A. ∎
Theorem 1. Based on Lemma 1, the angular-delay domain channel matrix $\mathbf{X}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}\mathbf{F}_{\mathrm{D}}^{*}$ exhibits a sparse structure, containing only $L$ nonzero blocks, each corresponding to an individual propagation path. For the $l$-th path, the support region of the nonzero components is denoted by $\mathcal{B}_{l}$, which can be expressed as $\mathcal{B}_{l}=\bigcup_{p}\mathcal{B}_{l,p}$, where $\mathcal{B}_{l,p}$ represents the shifted version of $\mathcal{A}_{l}$ in both the angular and delay domains, and is defined as
$\mathcal{B}_{l,p}$ =$=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l}+N\epsilon_{l,p},N),\right.$ (28)
$k$ $\left.=\mathrm{mod}(k_{l}+f_{s}\tau_{l},K),\forall(n_{l},k_{l})\in\mathcal{A}_{l}\right\},$
where $\mathrm{mod}(a,m)$ is the modulus of $a$ for $m$.
According to Theorem 1, the nonzero region of $\mathbf{X}$ is formed by aggregating the shifted nonzero regions of $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\psi_{l},\varphi_{l})\mathbf{F}_{\mathrm{D}}^{*}$ across both the angular and delay domains for all paths $l$. The shift positions are jointly determined by the spatial shift $N\epsilon_{l,p}$ and the delay shift $f_{s}\tau_{l}$, corresponding to the spatial frequency and propagation delay of each path, respectively.
Proof.
Please see Appendix B. ∎
Moreover, we analyze the effects of the spherical wavefront, SnS properties, and dual-wideband effects on the sparsity structure of $\mathbf{X}$ based on Lemma 1 and Theorem 1 from the perspectives of the angular and delay spread.
  • Angular Spread: First, the spherical wavefront effect and the SnS characteristics determine the angular spread range at each subcarrier, characterized by $i_{l,s}(k)$ and $i_{l,e}(k)$. Then, the frequency-wideband effect governs the overall boundaries of the angular spread, denoted by $I_{l,e}$ and $I_{l,s}$. Moreover, the frequency-wideband effect also enables a more accurate description of the angular spread within each column, represented by $B_{l,k}$.
  • Delay Spread: The spatial-wideband effect causes variations in $j_{l,e}(n)$ and $j_{l,s}(n)$ across antennas $n$, leading to delay spread. The SnS properties define the range of variation for $n$, while the spherical wavefront imposes a quadratic trend on the delay spread as a function of $n$.
@@
In summary, the spherical wavefront effect, SnS properties, and dual-wideband effects collectively govern the angular and delay spreads, thereby reshaping the sparsity structure of SnS dual-wideband channels in the angular-delay domain. Specifically,
@@
  • Global Block Sparsity: The angular-delay domain channel exhibits $L$ nonzero blocks, each corresponding to a distinct propagation path. The index range of each nonzero block is characterized by $\mathcal{B}_{l}$.
  • Local Common-Delay Sparsity: Due to the frequency-wideband effect, the significant components within $\mathcal{B}_{l}$ are confined to narrow bands along each column, thereby exhibiting pronounced column-wise clustered structures with strong correlation.
@@
(a) Angular spread for $k=0$ and $K-1$
(b) Delay spread for $n=n_{s}$ and $n_{e}$
(c) $\mathcal{A}_{l}$.
(d) $\mathcal{B}_{l,1}$.
(e) $\mathcal{B}_{l,P}$
(f) $\mathcal{B}_{l}$.
@@
To enhance intuitive understanding of the SnS XL-MIMO channel, we provide an example with the following parameters: $N=K=1024$, $f_{s}=3$GHz, $f_{c}=30$GHz, $r_{l}=10$m, $\vartheta_{l}=\pi/5$, and $\boldsymbol{\phi}_{l}\in[64,959]$. Fig. 2 illustrates the sparsity pattern of the SnS XL-MIMO channel in the angular-delay domain. Given the specified parameters, we have $i_{l,e}(0)=0.64$, $i_{l,s}(0)\approx 0.8$, $j_{l,s}(n_{s})=513$, and $j_{l,e}(n_{s})\approx 517$. Similarly, $i_{l,s}(K-1)\approx 0.88$, $i_{l,e}(K-1)\approx 0.72$, $j_{l,s}(n_{e})=545$, and $j_{l,e}(n_{e})\approx 549$. As depicted in Fig. 2(a) and Fig. 2(b), these values provide accurate boundaries for the ranges of angular and delay spread. Additionally, the figures show that both the angular and delay spectra shift to the right due to $\cos\vartheta_{l}>0$ and the monotonicity of $t_{n}$ with respect to $n$.
Furthermore, we obtain the following values: $I_{l,e}=841$, $I_{l,s}=962$, $J_{l,e}=513$, $J_{l,s}=549$, and $B_{k,s}\approx 41$ with $k=J_{l,e}$ Fig. 2(c) illustrates the non-zero region of $\mathbf{F}_{\mathrm{A}}^{{}_{\mathrm{H}}}\boldsymbol{\Theta}(\psi_{l},\varphi_{l})\mathbf{F}^{*}_{\mathrm{D}}$, which is accurately constrained to $\mathcal{A}_{l}$, validating the effectiveness of Lemma 1. In particular, for the $J_{l,e}$-th column, the width of angular spread also aligns the result of (27), which further validates the influence of frequency-wideband effect for local angular spread.
Additionally, we have $f_{s}\tau_{l}=100$, $\epsilon_{l,1}=\psi_{l}-2n_{l,s}\varphi_{l}\approx 0.4$, and $\epsilon_{l,P}=\psi_{l}-2n_{l,e}\varphi_{l}=0.32$, corresponding to $N\epsilon_{l,1}=410$ and $N\epsilon_{l,P}=328$, respectively. Consequently, Fig. 2(d) and Fig. 2(e) represent shifted versions of Fig. 2(c). The shift distances in the angular domain are $696$ and $614$, respectively, while the shift distance in the delay domain is $100$. Finally, by concentrating the shifted versions across all $\epsilon_{l,p}$, the sparse structure depicted in Fig. 2(f) is obtained.
@@
Compared to conventional massive MIMO-OFDM channels, the proposed SnS dual-wideband XL-MIMO channel exhibits distinct behavior. Specifically, in addition to beam squint, the spherical wavefront effect and SnS properties contribute to angular spread, resulting in broader angular diffusion. Moreover, the quadratic phase variations introduced by the spherical wavefront reshape the trend of delay variations across antennas. Notably, Lemma 1 is applicable to conventional MIMO-OFDM channels. When the spherical wavefront simplifies to a plane wavefront and SnS properties are disregarded, both $\varphi_{l}$ and $\kappa_{l,n}$ can be omitted. Thus, we have $i_{l,e}(k)=i_{l,s}(k)$ and $t_{l,n}=nf_{s}\psi_{l}$. Consequently, the significant components are uniformly distributed within the non-zero region, as illustrated in [9].
IV Problem Formulation and Sparse Prior Modeling
In this section, leveraging the sparsity outlined in Theorem 1, we first formulate the SnS dual-wideband XL-MIMO channel estimation problem as a sparse signal recovery task. Subsequently, to account for the global block sparsity and local common-delay sparsity, we introduce a novel column-wise sparse prior model tailored to this structure.
@@ IV-A Problem Formulation
Motivated by the sparsity of SnS dual-wideband XL-MIMO channels in the angular-delay domain, the channel estimation problem can be formulated as a MMV-based sparse recovery problem. Utilizing the angular-delay representation in (14), the received signal model in (3) can be further written as
$\mathbf{Y}=\mathbf{W}\mathbf{F}_{\mathrm{A}}\mathbf{X}\mathbf{F}_{\mathrm{D}}^{\mathrm{T}}+{\mathbf{N}}=\boldsymbol{\Psi}\mathbf{X}\mathbf{F}_{\mathrm{D}}^{\mathrm{T}}+{\mathbf{N}},$ (29)
@@
$\tilde{\mathbf{Y}}={\mathbf{Y}}\mathbf{F}_{\mathrm{D}}^{*}=\boldsymbol{\Psi}\mathbf{X}+\tilde{\mathbf{N}},$ (30)
where $\tilde{\mathbf{Y}}={\mathbf{Y}}\mathbf{F}_{\mathrm{D}}^{*}\in\mathbb{C}^{M\times K}$ and $\tilde{\mathbf{N}}={\mathbf{N}}\mathbf{F}_{\mathrm{D}}^{*}\in\mathbb{C}^{M\times K}$ denotes the equivalent received pilot signal and noise matrix.
Remark 4.
While this work primarily focuses on the ULA configuration, the signal models in (29) and (30) are also applicable to the uniform planar array (UPA) configuration. Specifically, consider a BS equipped with a half-wavelength UPA comprising $N=N_{v}\times N_{h}$ antennas, where $N_{v}$ and $N_{h}$ denote the number of elements along the vertical and horizontal dimensions, respectively. In this case, the angular transformation matrix $\mathbf{F}_{\mathrm{A}}$ can be redefined as the Kronecker product $\mathbf{F}_{\mathrm{A}}=\mathbf{F}_{\mathrm{A},v}\otimes\mathbf{F}_{\mathrm{A},h}$, where $\mathbf{F}_{\mathrm{A},v}\in\mathbb{C}^{N_{v}\times N_{v}}$ and $\mathbf{F}_{\mathrm{A},h}\in\mathbb{C}^{N_{h}\times N_{h}}$ denote the vertical and horizontal DFT matrices, respectively. It is worth noting that the sparsity structure in the angular-delay domain may differ between ULA and UPA configurations, particularly with regard to local common-delay sparsity, due to the inherent differences in array geometry and two-dimensional angular resolution. Although a comprehensive analysis of the UPA case is beyond the scope of this paper, we acknowledge its importance and leave it as a promising direction for future work.
In this paper, our primary objective is to develop an effective estimation algorithm for accurately reconstructing $\mathbf{X}$ or $\mathbf{H}$ according to $\mathbf{Y}$ and $\boldsymbol{\Psi}$. In the existing literature, several sparse signal recovery algorithms have been employed to address the problem in (30), including SOMP [30, 36] and optimization-based methods [37, 38]. However, these methods fail to effectively exploit the inherent sparsity structure of the angular-delay channel and often require additional parameters, such as the rank of $\mathbf{X}$ or the number of multipaths, which may not always be available. Considering the block sparsity of $\mathbf{X}$, Bayesian inference techniques such as approximate message passing (AMP) and SBL have shown superior recovery performance in channel estimation tasks [39, 40, 41]. However, the effectiveness of these Bayesian methods relies heavily on the accuracy of the prior model.